OpenStudy (anonymous):

Is this equivalent to 3?

OpenStudy (anonymous):

\[3x + 33\div x + 11\]

OpenStudy (anonymous):

Please explain so I understand.

OpenStudy (anonymous):

Is it \[\frac{3x + 33}{x+11}\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

If you factor 3 from the numerator, do you see something you can cancel out?

OpenStudy (anonymous):

@meepi I'm having a hard time understanding what exactly you mean.. Understanding math is like understanding Chinese to me>_<

OpenStudy (phi):

there are a few things you should learn if you have 3(x+11) think of the parens as a package, and 3(x+11) means you have 3 packages. that means you have 3 of each thing inside : 3 x's and 3 11's you write that as 3x+3*11 or 3x + 33

OpenStudy (phi):

factor out 3 undoes the 3*x + 3*11 you "factor out 3" to get 3(x+11) which is another way to right 3x+33

OpenStudy (phi):

another thing you should learn: anything divided by itself is 1 that works for anything *except zero* 2/2 = 1 3/3 = 1 (x+y+z)/(x+y+z) = 1 as long as it is not 0/0

OpenStudy (phi):

now to the question \[ \frac{3x + 33}{x+11} = \frac{3(x+11)}{(x+11)} \] as long as x+11 is not zero, we can divide (x+11)/(x+11) and get 1 and 3*1 is 3 so we can say \[ \frac{3x + 33}{x+11} = 3 \text{ but x≠ -11} \] so it is almost the same as 3... it is 3 except for when x=-11 when x= -11 you have 0/0 which is "indeterminate"... we don't know what it is.

OpenStudy (anonymous):

You. Are. Amazing.